OpenStudy (anonymous):

help medal show work 2.Once a commercial plane reaches the desired altitude,the pilot often travels at a cruising speed.On average,the cruising speed is 570 miles/hour.If s plane travels at this cruising speed for for 7 hours,how far does the plane travel while cruising at this speed.

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

could u show work

OpenStudy (anonymous):

how to multiply? i would use a calculator

OpenStudy (anonymous):

or draw it out

OpenStudy (anonymous):

no the work i ment

OpenStudy (anonymous):

do you know how to multiply \[570\times 7\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then that is the "work"

OpenStudy (anonymous):

k wait

OpenStudy (anonymous):

3909

OpenStudy (anonymous):

but there should be work not that much

OpenStudy (anonymous):

i have no idea how you got that number but i am certain that \[570\times 7\] is not that for one thing \[570\times 7\] ends in a zero

OpenStudy (anonymous):

3990

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=570*7

OpenStudy (anonymous):

yeah that looks better

OpenStudy (anonymous):

thanks for the help